OpenStudy (anonymous):

find a value for b so that f is a continuous function for all real numbers. Where f is defined f(x){ x^2 + bx , for x < or = 5 {5sin(pi/2(x)) , for x >5 do i use deriviative here?

OpenStudy (anonymous):

replace \(x\) by 5 in both expressions, set them equal and solve for \(b\) no, you do not need derivatives

OpenStudy (anonymous):

\[25+5b=5\sin(\frac{5\pi}{2})\] is a stsart

OpenStudy (anonymous):

so would b = -4?

OpenStudy (anonymous):

i think -5

OpenStudy (anonymous):

25+5b = 5sin(pi/2(5)) 25 + 5b = 5 5b = -20 b= -4?